The difference between the indices of carbon bi sulphide for blue light and red light is 0.48 while the critical angle for the red light at carbon bi sulphide air interface is 49°. Calculate the critical Ray of the carbon bi sulphide for the blue light. Thank you


1
Expert's answer
2021-02-11T10:41:02-0500

Instead of always having to measure the critical angles of different materials, it is possible to calculate the critical angle at the surface between two media using Snell's Law. To recap, Snell's Law states:

"n_1sin\u03b8_1=n_2sin\u03b8_2"

where "n_1" is the refractive index of material 1, "n_2"  is the refractive index of material 2,

"\u03b8_1"  is the angle of incidence and "\u03b8_2" is the angle of refraction. For total internal reflection, we know that the angle of incidence is the critical angle. So,

"\u03b8_1=\u03b8 _c."

However, we also know that the angle of refraction at the critical angle is 90°. So we have:

"\u03b8_2=90^\u2218."

We can then write Snell's Law as:

"n_1sin\u03b8_c=n_2sin90^\u2218"

Solving for "\u03b8_c"  gives:

"n_1sin\u03b8\n_c = n_2sin90\u00b0"

"sinθ_c = \large\frac{n_2}{n_1}"

"∴θ_c=sin^{−1}(\large\frac{n_2}{n_1})"

We have "n_2 = 1" for air "\theta_c = 48^o"

"sin48^o = \large\frac{1}{n_1} \to 0.74n_1 = 1 \to n_1 = 1.35"

"n_2 = n_1 + 0.48 = 1.81"

Critical ray for carbon bi sulphide for the blue light

"sin\theta _c = \frac{n_2}{n_1} = \frac{1}{1.85} = 0.54 \to \theta_c = 33.5^o"